Circular No. 7470

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000cz IN IC 1535 G. M. Hurst, Basingstoke, England, reports the discovery by M. Arm-

strong, Rolvenden, Kent, of an apparent supernova (mag 18) on four CCD patrol images (limiting mag 19–19.5) taken with a 0.30-m reflector in the course of the U.K. Nova/Supernova Patrol on July 29.048–29.108 UT. From a 60-s image obtained on July 29.981, in which the new object still appears at mag 18 (limiting mag 19.5), Armstrong reports the following precise position for SN 2000cz: α = 0^{h}13^{m}58^{s}.19, δ = +48^{o}09′14′′.4 (equinox 2000.0). Nothing appears at this position on four previous images of this galaxy taken by Armstrong on 1999 Nov. 2 (limiting mag 19), and nothing appears on second Palomar Sky Survey red (limiting mag 20.8) and blue (limiting mag 22.5) plates. Hurst adds that R. Arbour, South Wonston, Hampshire, obtained a prediscovery image of SN 2000cz on July 21.9.

W. D. Li, University of California at Berkeley, reports that SN 2000cz also appeared in images taken by the Katzman Automatic Imaging Telescope on July 8.5 (mag ∼ 17.7), 17.5 (mag ∼ 17.9), and 27.5 UT (mag ∼ 18.0). From the image on July 27.5, Li also measured the following position end figures for SN 2000cz: 58^{s}.23, 14′′.2, which is 8′′.4 east and 14′′.9 south of the nucleus of IC 1535.

COMET C/1999 S4 (LINEAR) Further to their item on IAUC 7468, Filippenko and Chornock report

that full reduction of their CCD spectrum obtained on July 28 shows that CN emission near 380 nm is actually present; the CN was not initially noticed in the two-dimensional spectrum because it spans the entire slit, making it resemble a night-sky emission line. In contrast, the continuum (reflected sunlight) is visible only on part of the slit, ending abruptly (as mentioned on IAUC 7468).

SUPERNOVA 2000cx IN NGC 524 V. Vlasyuk and N. Borisov, Special Astrophysical Observatory, report

that spectra (range 390–790 nm; resolution ∼ 1.2 nm; signal-to-noise ratio 50) of SN 2000cx (cf. IAUC 7458, 7463), obtained over 60 min of exposure time with the 1-m reflector (+ long-slit spectrograph + CCD camera) on July 26.99 UT, reveal spectral features and a spectral slope that are typical for type-Ia supernovae, and in particular very similar to that of SN 1998aq near maximum.

2000 July 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green